If the complement of an angle is 10° less than one-third of its supplement, find the measure of the angle.
i have tried .3(180)-10=90-x and webassign.net says it is not 40, and it is not 50 either.

Answer by stanbon(75887)

If the complement of an angle is 10° less than one-third of its supplement, find the measure of the angle.
---
Let the angle be "x".
Its complement is 90-x
Its supplement is 180-x
----
Equation:
Solve for "x":
90-x = (1/3)(180-x)-10
Multiply thru by 3 to get rid of the fractions:
270-3x = 180-x - 30
270-3x = 150-x
2x = 120
x = 60
============
Cheers,
Stan H.


i have tried .3(180)-10=90-x and webassign.net says it is not 40, and it is not 50 either.